a) Express the equation in the form y = mx + b,
b) find the slope and the y-intercept and
c) sketch the graph of the linear function:
y - 2 = -3 (x-1)

Answer by mananth(16946)

y - 2 = -3 (x-1)
y-2 =-3x+3
y=-3x+5
slope m = -3 and y intercept = 5